• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M. Rvachev

    Log entry time 02:01:15 on December 21,1999

    Entry number 31639

    keyword=EQ3 quench

    EQ3 quenched for no apparent reason (except that it was at 3.9GeV).
    Went down in the hall and reset it. The resetting procedure is pretty simple:

    1. Go to the quad power supply control boxes in the rear of spectrometer (use a movable ladder).
    2. Find the right box (look for red "quench" and "interlock" leds on).
    3. Put the big power handle on the box down (off).
    4. Press two "reset" buttons on the box panel.
    5. Wait ~30sec until red "interlock" and "quench" leds go off.
    6. Turn the big power hadle up (on).